OpenStudy (anonymous):

Part 1: Create a unique perfect square trinomial using what you know about the characteristics of a perfect square trinomial. Explain why the trinomial you created is considered a perfect square trinomial.

OpenStudy (anonymous):

It has to do with completing a square

OpenStudy (anonymous):

if you square a binomial, you end up with a "perfect square trinomial":\[(x+y)^2 = x^2+2xy+y^2\]So just pick values for y to come up with some trinomials. if I pick y = 3 i get: \[(x+3)^2 = x^2+6x+9\]thats a perfect square trinomial. Just pick any number for y.

OpenStudy (anonymous):

Simplisticly,When the third term is halfed and squared of the second and first term has 1 as coefficient .

OpenStudy (anonymous):

what if i did this: (x+2)(x+4) x^2+8+2x+4x x^2+6x+8 would that be right??

OpenStudy (anonymous):

thats not a perfect square though. If you had done:\[(x+2)(x+2) = (x+2)^2\]then you would be fine.

OpenStudy (anonymous):

i dont understand. they are equal

OpenStudy (anonymous):

you want two factor to be the same ex, (x+2)(x+2) or (x-3)(x-3) (x-a)(x-a)

OpenStudy (anonymous):

oh!! I see :P thank you guys so much

OpenStudy (anonymous):

how would i explain why it is a perfect square trinomial

OpenStudy (mathmate):

A TRInomial is typically of the form (a+b+c) So a perfect square for a trinomial is (a+b+c)^2 =a^2+b^2+c^2+2ab+2bc+2ca

OpenStudy (anonymous):

1) x^2+4x+4 This is a perfect square trinomial because first and third terms are perfect squares. 2) Find two numbers that multiply and add up to 4. The numbers are 2&2 (x+2)(x+2)=(x+2)^2

OpenStudy (anonymous):

hmm...i guess we are interpreting what this means in different ways.

OpenStudy (mathmate):

I think your interpretation is more appropriate. Sorry.

OpenStudy (anonymous):

meh, it is a little ambiguous, no worries :)

OpenStudy (anonymous):

lol this is alg 1. I am confused on defferent parts of it :P such as factoring